### v1

# Note on supersymmetric mechanics with spin-orbit interaction

Abstract
We propose a simple model of two-dimensional $\mathcal{N}=2$ superconformal mechanics with a spin-orbit interaction term and demonstrate that it inherits the Galilean symmetry of the initial free-particle system. We then propose a quaternionic counterpart of this system, which describes the four-dimensional $\mathcal{N}=4$ superconformal mechanics $D(1,2|\alpha)$. The bosonic part of the Hamiltonian describes a free particle on the cone, while the fermionic part necessarily includes the spin-orbit interaction term.
I Introduction
Supersymmetric mechanics was initially introduced as toy models for supersymmetric field theories [1]. However, it was soon realized that these models are highly relevant to the study of numerous problems in conventional quantum mechanics (for a review, see [2] and references therein). Nevertheless, the superfield approach—inspired by supersymmetric field theories—remains the primary tool for studying of supersymmetric mechanics. Over the past two decades, significant efforts have been devoted to constructing models of $\mathcal{N}\geq 4$ supersymmetric mechanics within this framework. In particular, all $\mathcal{N}=8$ were classified in [3]. Much attention has been devoted to superconformal mechanics (see, e.g., [4] and references therein).
Many of these models were formulated using Lagrangians that depend on additional bosonic variables introduced via first-order kinetic terms (see, e.g., [5]). These variables were interpreted by the authors as isospin (or semidynamical) variables. Recently, Fedoruk and Ivanov proposed a highly nontrivial $\mathcal{N}=8$ supersymmetric mechanical model of this kind [6]. The model was formulated in terms of a Lagrangian depending on a single real bosonic variable, eight complex bosonic variables, and eight real fermionic variables. Only the real bosonic variable appears with a second-order kinetic term, while the kinetic terms for the complex bosonic variables are of first order. The authors interpreted the model as an $\mathcal{N}=8$ supersymmetric extension of a one-dimensional isospin particle.
Shortly afterward, together with Erik Khastyan, we demonstrated that the Fedoruk-Ivanov model possesses dynamical superconformal symmetry and describes an $\mathcal{N}=8$ supersymmetric extension of a free particle on an eight-dimensional cone. Additionally, its fermionic part can be interpreted as a spin-orbit coupling term [7]. In [8], this construction was extended to a broader class of $\mathcal{N}=8$ superconformal mechanics. We also argued that the Fedoruk-Ivanov model has an intrinsic connection to octonions, suggesting that similar $\mathcal{N}=2,4$ models exist in two- and four-dimensional spaces and are associated with complex numbers and quaternions, respectively. In this note, we demonstrate that this is indeed the case.
In Section 2, we propose and describe in detail an $\mathcal{N}=2$ supersymmetric mechanics model with two bosonic and two fermionic degrees of freedom. The bosonic part of the Hamiltonian corresponds to a free particle on a two-dimensional plane, while the fermionic part defines a spin-orbit interaction term. Moreover, the system inherits the Galilean symmetry of the two-dimensional free particle and possesses the dynamical $su(1,1|1)$ superconformal symmetry.
In Section 3, we present the quaternionic counterparts of the supercharges introduced in the previous section and construct, in this way, an $\mathcal{N}=4$ supersymmetric mechanics model with four bosonic and four fermionic degrees of freedom. This system exhibits the dynamical superconformal symmetry $D(1,2|\alpha)$ and is formulated on a four-dimensional Euclidean cone, which reduces to $\mathbf{R}^{4}$ for $\alpha=1$. The fermionic part of the Hamiltonian necessarily includes a spin-orbit coupling term.
II $\mathcal{N}=2$ supersymmetric systems
Consider the two-dimensional Euclidean plane ${\mathbf{{R}}^{2}}={\mathbf{C}}^{1}$ equipped with the metrics
$(ds)^{2}=\sum_{\mu=0,1}dx^{\mu}dx^{\mu}=2dzd\bar{z},\qquad{\rm with}\quad z=\frac{x^{0}+\imath x^{1}}{\sqrt{2}}.$ (1)
Then extend this space by two fermionic degrees of freedom $\theta^{1},\theta^{2}$ and introduce the momenta $p_{0},p_{1}$ canonically conjugated with $x^{0},x^{1}$. This extended superspace can be equipped with the canonical (super)symplectic structure
$\Omega=\sum_{\mu=0,1}\left(dp_{\mu}\wedge d{x}^{\mu}+\frac{1}{2}d\theta^{\mu}\wedge d\theta^{\mu}\right)=d\pi\wedge dz+d\bar{\pi}\wedge d\bar{z}+d\eta\wedge d\bar{\eta},$ (2)
where
$z=\frac{x^{0}+\imath x^{1}}{\sqrt{2}},\qquad\pi=\frac{p_{0}-\imath p_{1}}{\sqrt{2}},\qquad\eta=\frac{\theta^{0}+\imath\theta^{1}}{\sqrt{2}}.$ (3)
The respective Poisson brackets are given by the non-zero relations
$\{\pi,z\}=1,\quad\{\eta,\bar{\eta}\}=1\quad\Leftrightarrow\quad\{p_{\mu},x^{\nu}\}=\delta_{\mu}^{\nu},\quad\{\theta^{\mu},\theta^{\nu}\}=\delta^{\mu\nu}.$ (4)
Our goal is to construct, on this space, the $\mathcal{N}=2$ supersymmetric extension of the free particle system defined by the Hamiltonian $H=\pi\bar{\pi}$.
For this purpose we choose the following Ansatz for supercharge:
$Q={\rm e}^{\imath\nu\;arg\;z}\pi\eta,\qquad{\rm with}\quad{\rm e}^{\imath\;arg\;z}=\frac{z}{|z|},\quad\nu\in\mathcal{N}.$ (5)
It immediately yields the $\mathcal{N}=2$ supersymmetric Hamiltonian
$\mathcal{H}=\pi\bar{\pi}+\frac{\nu}{2}\frac{(\pi z-\bar{\pi}\bar{z})\eta\bar{\eta}}{|z|^{2}}\;:\qquad\{Q,\overline{Q}\}=\mathcal{H},\quad\{Q,Q\}=0.$ (6)
The Hamiltonian and supercharges commute with the following constants of motion:
${{\Pi}}=\pi-\frac{\nu}{2z}\eta\bar{\eta}\;,\quad\mathcal{J}=\imath(\pi z-\bar{\pi}\bar{z})+\imath(1-\nu)\eta\bar{\eta}:$ (7)
$\{{\Pi},\mathcal{H}\}=\{\mathcal{J},\mathcal{H}\}=0,\quad\{{\Pi},Q\}=\{{\Pi},\bar{Q}\}=\{\mathcal{J},Q\}=0$ (8)
The algebra of these constants of motion is of the form
$\{{\Pi},{\bar{\Pi}}\}\approx 0,\quad\{{\Pi},\mathcal{J}\}=\imath\Pi,$ (9)
i.e. the constructed system inherits symmetries of the two-dimensional free particle [^1] The notation $\approx$ in (9) is not a misprint because to accurately calculate the Poisson brackets between $\Pi,\bar{\Pi}$ we should have in mind the relation $\frac{\partial}{\partial z}\frac{1}{\bar{z}}=2\pi\imath$ so that $\{\Pi,\bar{\Pi}\}=2\pi\nu\imath\delta^{2}(z)\imath\eta\bar{\eta}$ ( here $\pi$ is not a momentum but the irrational number: $\pi=3,1415...$ ). After quantization the operator $\Pi$ acts on the upper/lower component of the two-component spinor as $\partial_{z}\mp\imath\nu/z$. Since $A_{z}=\frac{\imath}{2\mathcal{\pi}z}$ is the vector potential of the infinitely thin magnetic solenoid (in two-dimensional terms - magnetic vortex), the system can be interpreted as an anyon with spin $\nu$, or, equivalently, as a charged particle moving in the field of an infinitely thin solenoid, see, e.g. [9]..
Performing canonical transformation mixing bosonic and fermionic degrees of freedom, one could represent the system in the form of a free particle. Namely, introducing the new fermionic variable
$\tilde{\eta}={\rm e}^{\imath\nu\;arg\;z}\eta\;,$ (10)
we get
$\{\tilde{\eta},\bar{\tilde{\eta}}\}=1,\quad\{\Pi,z\}=1,\quad\{\bar{\Pi},\bar{z}\}=1,\qquad\{\Pi,\tilde{\eta}\}=\{\Pi,\bar{\tilde{\eta}}\}=0.$ (11)
In these terms the Hamiltonian, supercharges, and rotation generator read
$\mathcal{H}=\Pi\bar{\Pi},\quad Q=\Pi\tilde{\eta},\quad\overline{Q}={\bar{\Pi}}\bar{\tilde{\eta}},\quad\mathcal{J}=\imath(\Pi z-\bar{\Pi}\bar{z})+\imath{\tilde{\eta}}\bar{\tilde{\eta}}\;.$ (12)
The above $\mathcal{N}=2$ supersymmetry algebra can be immediately extended to the superconformal one by adding the generators
$K={z\bar{z}},\quad D=z\pi+\bar{z}\bar{\pi}=z\Pi+\bar{z}\bar{\Pi},\quad\tilde{\mathcal{J}}=\mathcal{J}+\imath\eta\bar{\eta},\quad S={\rm e}^{\imath\nu\;arg\;z}\bar{z}\eta=\bar{z}\tilde{\eta}:$ (13)
$\{K,\mathcal{H}\}=-D,\quad\{D,K\}=2K,\quad\{D,\mathcal{H}\}=-2\mathcal{H},$ (14)
$\{D,Q\}=-Q,\quad\{D,S\}=S,\quad\{K,Q\}=-S,\quad\{S,Q\}=0,$ (15)
$\{S,\bar{Q}\}=\frac{1}{2}\left(D+\imath\tilde{\mathcal{J}}\right),\quad\{\mathcal{H},S\}=Q,\quad\{S,\bar{S}\}=K,\quad\{\tilde{\mathcal{J}},Q\}=\imath Q.$ (16)
Quantization of similar two-dimensional model (with additional oscillator term and broken $\mathcal{N}=2$ supersymmetry) was given in [10].
Let us complete this section by two remarks:
  • The bosonic and fermionic parts of the rotational momentum $\mathcal{J}$ (and of $\tilde{\mathcal{J}}$ ) define the constants of motion as well, but they do not commute with supercharges
    $L:=\imath(\pi z-\bar{\pi}\bar{z}),\quad R:=\imath\eta\bar{\eta}\;:$ (17)
    $\quad\{L,\mathcal{H}\}=\{R,\mathcal{H}\}=\{L,R\}=0,\quad\{L,Q\}=\imath(\nu-1)Q,\quad\{R,Q\}=\imath Q.$ (18)
    In the three-dimensional terms $L$ corresponds to the third component of the angular momentum while $R$ upon quantization results in the Pauli matrix $\sigma_{3}$. So, the fermionic part of the Hamiltonian defines the spin-orbit coupling term. It is the well-known system describing the so-called spin-Hall effect observed two decades ago [11]. The supercharge is nothing else but the Hamiltonian of the Rashba model.To our knowledge, a supersymmetric features of that model have not been noticed in the extended literature on this subject.
  • The constructed supersymmetric system preserves all symmetry properties upon the inclusion of the constant magnetic field. This can be done by an appropriate deformation of the symplectic structure
    $\Omega_{B}=\Omega+\imath Bdz\wedge d\bar{z}$ (19)
    and of the respective Poisson brackets
    $\{\pi,z\}_{B}=1,\quad\{\eta,\bar{\eta}\}_{B}=1,\quad\{\pi,\bar{\pi}\}_{B}=\imath B,$ (20)
    where $B=const$ is the strength of the magnetic field.
    Then, choosing the supercharge in the form (5), we arrive at the Hamiltonian
    $\mathcal{H}_{B}:=\{Q,\overline{Q}\}_{B}=\mathcal{H}+\imath B\eta\bar{\eta}.$ (21)
    In this case, the translation and rotation generators take the form
    $\Pi_{B}=\Pi+\imath B\bar{z},\qquad\mathcal{J}_{B}=\mathcal{J}-{Bz\bar{z}}.$ (22)
    They remain commuting with the Hamiltonian and supercharges and form the two-dimensional Galilieo algebra.
Let us rewrite the above expressions in the polar coordinates
$z=\frac{r{\rm e}^{\imath\nu\varphi}}{\sqrt{2}},\quad\pi=\frac{{\rm e}^{-\imath\nu\varphi}}{\sqrt{2}}\left(p_{r}-\frac{\imath}{\nu}\frac{p_{\varphi}}{r}\right)\;:\quad\{p_{r},r\}=1,\quad\{p_{\varphi},\varphi\}=1.$ (23)
In these terms the Hamiltonian and supercharges take the following form
$\mathcal{H}=\frac{p^{2}_{r}}{2}+\frac{p^{2}_{\varphi}}{2\nu^{2}r^{2}}+\frac{p_{\varphi}\theta^{1}\theta^{2}}{2r^{2}},\qquad Q=\frac{1}{\sqrt{2}}\left(Q_{1}+\imath Q_{2}\right),\qquad Q_{1}=p_{r}\theta^{1}+\frac{p_{\varphi}\theta^{2}}{\nu r},\quad Q_{2}=p_{r}\theta^{2}-\frac{p_{\varphi}\theta^{1}}{\nu r}\;.$ (24)
The generators of the Galilieo algebra read
$\mathcal{J}=\frac{p_{\varphi}}{\nu}+(1-\nu)\theta^{1}\theta^{2},\quad\Pi=\Pi_{1}+\imath\Pi_{2}=\frac{{\rm e}^{-\imath\nu\varphi}}{\sqrt{2}}\left(p_{r}-\imath\frac{p_{\varphi}-\nu\theta^{1}\theta^{2}}{r}\right)$ (25)
These expressions will be used in the next section to choose the Ansatz for the $\mathcal{N}=4$ supersymmetric mechanics model.
III $\mathcal{N}=4$ supersymmetric system
It this Section we propose the quaternionic counterpart of the $\mathcal{N}=2$ supersymmetric systems constructed in the previous Section. We expect to obtain in this way the $\mathcal{N}=4$ supersymmetric extensions of the free particle systems on the four-dimensional Euclidean space $\mathbf{{R}}^{4}=\mathbf{{H}}^{1}$ equipped with the metrics
$(ds)^{2}=\sum_{\mu=0}^{3}dx^{\mu}dx^{\mu}=2d\mathbf{z}d\bar{\mathbf{z}},\qquad z=\frac{x^{0}+\sum_{i=1}^{3}\mathbf{e}_{i}x^{i}}{\sqrt{2}},$ (26)
where
$\mathbf{e}_{i}\mathbf{e}_{j}=-\delta_{ij}+\epsilon_{ijk}\mathbf{e}_{k},$ (27)
and $\epsilon_{ijk}$ is the completely antisymmetric tensor with $\epsilon_{123}=1$.
Then we extend this space by four fermionic degrees of freedom defined by the coordinates $\theta^{0},\theta^{i}$ and introduce the momenta $p_{0},p_{i}$ canonically conjugated with $x^{0},x^{i}$.
This extended superspace can be equipped with the (super)canonical Poisson brackets given by the following nonzero relations:
$\{p_{\mu},x^{\nu}\}=\delta_{\mu}^{\nu},\qquad\{\theta^{\mu},\theta^{\nu}\}=\delta^{\mu\nu}.$ (28)
The respective (super)symplectic structure can be re-written in quaternionic coordinates in a form similar to (2)
$\Omega=\mathbf{Re}\;\left(d\boldsymbol{\pi}\wedge d\mathbf{z}+d\boldsymbol{\eta}\wedge d\boldsymbol{\bar{\eta}}\right)=\sum_{\mu=0}^{3}\left(dp_{\mu}\wedge d{x}^{\mu}+\frac{1}{2}d\theta^{\mu}\wedge d\theta^{\mu}\right),$ (29)
where
$\boldsymbol{\pi}=\frac{p_{0}-\sum_{i=1}^{3}p_{i}\mathbf{e}_{i}}{\sqrt{2}},\quad\boldsymbol{z}=\frac{x^{0}+\sum_{i=1}^{3}x^{i}\mathbf{e}_{i}}{\sqrt{2}},\quad\boldsymbol{\eta}=\frac{\theta^{0}+\sum_{i=1}^{3}\theta^{i}\mathbf{e}_{i}}{\sqrt{2}}.$ (30)
Before extending the Anzats (5) to the quaternionic case, we present some quaternionic relations that will be further used for the construction of $\mathcal{N}=4$ supersymmetric mechanics.
At first, let us write down the relations on bosonic variables
$\frac{\boldsymbol{z\pi}}{|\mathbf{z}|}=\frac{1}{\sqrt{2}}\left(p_{r}+\frac{L^{+}_{i}\mathbf{e}_{i}}{r}\right),\qquad\frac{\boldsymbol{\pi z}}{|\mathbf{z}|}=\frac{1}{\sqrt{2}}\left(p_{r}+\frac{L^{-}_{i}\mathbf{e}_{i}}{r}\right),$ (31)
where
$r:=\sqrt{2}|\mathbf{z}|=\sqrt{\sum_{\mu=0}^{3}x^{\mu}x^{\mu}},\quad p_{r}=\sqrt{2}\frac{\mathbf{Re}\;\boldsymbol{z\pi}}{|z|}=\frac{\sum_{\mu=0}^{3}p_{\mu}x^{\mu}}{r},$ (32)
and
$L^{\pm}_{i}=p_{0}x^{i}-p_{i}x^{0}\pm\epsilon_{ijk}p_{j}x^{k}\;:\{L^{\pm}_{i},L^{\pm}_{j}\}=\pm 2\epsilon_{ijk}L^{\pm}_{k},\quad\{L^{+}_{i},L^{-}_{j}\}=0.$ (33)
So $L^{\pm}_{i}$ are the generators of the $so(3)$ algebras corresponding to the decomposition $so(4)=so(3)\times so(3)$ with the Casimirs
$\mathcal{C}_{so(3)}=\sum_{i=1}^{3}(L^{\pm}_{i})^{2}=\left(\sum_{\mu=0}^{3}p^{2}_{\mu}\right)\left(\sum_{\mu=0}^{3}(x^{\mu})^{2}\right)-\left(\sum_{\mu=0}^{3}p_{\mu}x^{\mu}\right)^{2}.$ (34)
Hence $\mathcal{C}_{so(4)}=2\mathcal{C}_{so(3)}$.
The Poisson brackets between the functions (32) and the $so(3)$ generators (33) are as follows:
$\{p_{r},r\}=1,\qquad\{L^{\pm}_{i},r\}=\{L^{\pm}_{i},p_{r}\}=0.$ (35)
Let also introduce the fermionic functions
$\mathbf{R}^{+}=\boldsymbol{\bar{\eta}}\boldsymbol{\eta}=R^{+}_{i}\mathbf{e}_{i},\quad\mathbf{R}^{-}=\boldsymbol{\eta}\boldsymbol{\bar{\eta}}=R^{-}_{i}\mathbf{e}_{i},\quad\mathbf{R}^{0}=\boldsymbol{\eta}^{2}=\boldsymbol{\bar{\eta}}^{2}=\frac{1}{2}R^{0}_{i}\mathbf{e}_{i}$ (36)
with
$R^{\pm}_{i}=\theta^{0}\theta^{i}\mp\frac{1}{2}\epsilon_{ijk}\theta^{j}\theta^{k},\quad R^{0}_{i}=R^{-}_{i}-R^{+}_{i}=\epsilon_{ijk}\theta^{j}\theta^{k}.$ (37)
They obey the following Poisson bracket relations:
$\{R^{\pm}_{i},R^{\pm}_{j}\}=\pm 2\epsilon_{ijk}R^{\pm}_{k},\quad\{R^{+}_{i},R^{-}_{j}\}=0,\quad\{R^{\pm}_{i},R^{0}_{j}\}=-2\epsilon_{ijk}R_{k}^{\pm},\quad\{R^{0}_{i},R^{0}_{j}\}=-2\epsilon_{ijk}R_{k}^{0}$ (38)
i.e., generate $so(4)$ (and $so(3)$) transformations in the fermionic sector.
It is instructive to write down the component expressions of the three-fermionic functions
$\boldsymbol{\Lambda}=\boldsymbol{\eta\bar{\eta}\eta}=\frac{\Lambda_{0}+\mathbf{e}_{i}\Lambda_{i}}{6\sqrt{2}}\;:\;\Lambda_{0}=\frac{1}{3}\epsilon_{ijk}\theta^{i}\theta^{j}\theta^{k},\quad\Lambda_{i}=-\epsilon_{ijk}\theta^{j}\theta^{k}\theta^{0},\qquad\{\Lambda_{0},\Lambda_{i}\}=\{\Lambda_{i},\Lambda_{j}\}=0$ (39)
With these expressions at hand we are ready to construct the $\mathcal{N}=4$ supersymmetric mechanics on $\mathbb{R}^{4}=\mathbb{H}^{1}$.
Due to noncommutativity of quaternions, there are many ways to extend the Ansatz (5) to the quaternionic case, including its extension by terms trilinear on fermions , e.g.,
$\boldsymbol{Q}=\left(\frac{\mathbf{z}}{|\mathbf{z}|}\right)^{\nu_{1}}\boldsymbol{\pi}\left(\frac{\mathbf{z}}{|\mathbf{z}|}\right)^{\nu_{2}}\boldsymbol{\eta}+\mu\left(\frac{\mathbf{z}}{|\mathbf{z}|}\right)^{\nu_{3}}\frac{\boldsymbol{\Lambda}}{|\mathbf{z}|}.\quad$ (40)
However, the absence of differentiation by quaternionic variables makes them not very effective for our purposes.
The simplest quaternionic analog of the $\mathcal{N}=2$ supercharge with $\nu=1$ (5) reads:
$\mathbf{\mathcal{Q}}=\frac{\mathbf{z}}{|\mathbf{z}|}\boldsymbol{\pi}\boldsymbol{\eta}=\frac{1}{\sqrt{2}}\left(\mathcal{Q}_{0}+\mathcal{Q}_{i}\mathbf{e}_{i}\right)\;:\qquad\mathcal{Q}_{0}=p_{r}\theta^{0}+\frac{L^{+}_{k}\theta^{k}}{r},\quad\mathcal{Q}_{i}=p_{r}\theta^{i}-\frac{L^{+}_{i}\theta^{0}-\epsilon_{ijk}L^{+}_{j}\theta^{k}}{r}$ (41)
However,one can check that these functions do not form the $\mathcal{N}=4$ superalgebra.
To get the $\mathcal{N}=4$ superalgebra, we choose a more general Ansatz including the three-fermionic term and depending on the coefficient $\alpha$ which could be viewed as the analog of the coefficient $\nu$ in (5)(cf. (24)).
Then, after direct calculation we get the expressions for supercharges which obeys $\mathcal{N}=4$ supersymmetry algebra
$Q_{0}=p_{r}\theta^{0}-\alpha\frac{L^{+}_{k}\theta^{k}}{r}+(1+2\alpha)\frac{\Lambda_{0}}{2r},\qquad Q_{i}=p_{r}\theta^{i}+\alpha\frac{L^{+}_{i}\theta^{0}-\epsilon_{ijk}L^{+}_{j}\theta^{k}}{r}+(1+2\alpha)\frac{\Lambda_{i}}{2r},$ (42)
Namely.
$\{Q_{0},Q_{0}\}=2\mathcal{H},\quad\{Q_{i},Q_{j}\}=2\delta_{ij}\mathcal{H},\qquad\{Q_{0},Q_{i}\}=0,$ (43)
where the Hamiltonian is given by the expression
$\mathcal{H}=\frac{1}{2}p_{r}^{2}+\frac{1}{2r^{2}}\left[{\alpha^{2}}{L^{+}_{k}L^{+}_{k}}+2\alpha{L^{+}_{k}R^{+}_{k}}-(1+2\alpha)\Gamma\right],\qquad{\rm with}\quad\Gamma:=\mp\frac{1}{3}{R^{\pm}_{k}R^{\pm}_{k}}=\frac{1}{3}{\theta^{0}\epsilon_{ijk}\theta^{i}\theta^{j}\theta^{k}}.$ (44)
It is an obvious that the following sets of $so(3)$ generators define the constants of motion of the Hamiltonian
$L^{-}_{i},\quad R^{-}_{i},\quad\mathcal{J}^{+}_{i}=L^{+}_{i}+R^{+}_{i}\;:\quad\{{L}^{-}_{i},\mathcal{H}\}=\{R^{-}_{i},\mathcal{H}\}=\{\mathcal{J}^{+}_{i},\mathcal{H}\}=0.$ (45)
Their Poisson brackets with supercharges are as follows:
$\{\mathcal{J}^{+}_{i},Q_{0}\}=-Q_{i},\quad\{\mathcal{J}^{+}_{i},Q_{j},\}=\delta_{ij}Q_{0}+\epsilon_{ijk}Q_{k},$ (46)
$\{{R}^{-}_{i},Q_{0}\}=-Q_{i},\quad\{{R}^{-}_{i},Q_{j}\}=\delta_{ij}Q_{0}-\epsilon_{ijk}Q_{k},$ (47)
$\{Q_{0},L^{-}_{k}\}=\{Q_{i},L^{-}_{k}\}=0,$ (48)
This superalgebra can be expended to the superconformal algebra $D(1,2|\alpha)$ by introducing the generators of dynamical superconformal symmetry
$K=\frac{r^{2}}{2},\quad D=\frac{rp_{r}}{2},\quad S_{0}=r\theta_{0},\quad S_{i}=r\theta_{i}.$ (49)
Their Poisson brackets with the former generators are as follows:
$\left\{S_{0},Q_{0}\right\}=2D,\quad\left\{S_{0},Q_{i}\right\}=\alpha{\cal J}^{+}_{i}-(1+\alpha)R^{-}_{i},$ (50)
$\left\{S_{i},Q_{0}\right\}=-\alpha{\cal J}^{+}_{i}+(1+\alpha)R^{-}_{i},\quad\left\{S_{i},Q_{j}\right\}=2\delta_{ij}D-\epsilon_{ijk}(\alpha{\cal J}^{+}_{k}+(1+\alpha)R^{-}_{k}),$ (51)
$\left\{\mathcal{J}^{+}_{i},S_{0}\right\}=-S_{i},\quad\left\{\mathcal{J}^{+}_{i},S_{j}\right\}=\delta_{ij}S_{0}+\epsilon_{ijk}S_{k},\quad\left\{R^{-}_{i},S_{0}\right\}=-S_{i},\quad\left\{R^{-}_{i},S_{j}\right\}=\delta_{ij}S_{0}-\epsilon_{ijk}S_{k}.$ (52)
The bosonic part of the Hamiltonian in the initial Cartesian coordinates reads
$H_{0}:=\frac{1}{2}p_{r}^{2}+\frac{{\alpha^{2}}L^{+}_{k}L^{+}_{k}}{2r^{2}}=\frac{\alpha^{2}}{2}\left(\sum_{\mu=0}^{3}p_{\mu}p_{\mu}-\frac{1-\alpha^{2}}{\alpha^{2}}\sum^{3}_{\mu=0}\left(\frac{p_{\mu}x^{\mu}}{|x|}\right)^{2}\right).$ (53)
So it describes the particle on the four-dimensional space equipped with the metrics
$(ds)^{2}=\sum_{\mu=0}^{3}dx^{\mu}dx^{\mu}-(1-\alpha^{2})\sum_{\mu=0}^{3}\left(\frac{x^{\mu}dx^{\mu}}{|x|}\right)^{2}=\frac{\sum_{\mu=0}^{3}dy^{\mu}dy^{\mu}}{|y|^{2(1-1/|\alpha|)}},$ (54)
where $y^{\mu}=|x|^{|\alpha|-1}x^{\mu}$.
Let us complete this section by the following remarks
  • For $\alpha=-1$, i.e. for the $su(1,1|2))$ superalgebra, we get the system on the four-dimensional Euclidian space, while for $\alpha^{2}\neq 1$ we deal with the free particle on four-dimensional cone.
  • For $\alpha=-1/2$, i.e., for the $osp(4|2)$ superalgebra, the four-fermionic (i.e spin-spin interaction) term vanishes and supersymmetry is provided solely by spin-orbit interaction.
  • The suggested model is a a particular example of $\mathcal{N}=4$ supersymmetric mechanics with the spin-orbit interaction term. For example, one can choose a more general Ansatz for supercharges, including both sets of $so(3)$ generators: not only $L^{+}_{\mu},R^{+}_{\mu}$ but $L^{-}_{\mu},R^{-}_{\mu}$ as well. We are planning to consider such models elsewhere.
  • We expect that a similar model could be constructed on the curved $so(4)$-invariant spaces by modifying the expressions for superchages and Hamiltonian presented in [12]. Clearly, in this case we will loose the dynamical superconformal symmetry.
Acknowledgements.
The work of A.N. was partially supported by the Armenian State Committee of Higher Education and Science within the project 21AG-1C062 .